Find the complete integral of
p²+q²-2px-2qy+1 =0
1
Expert's answer
2019-08-26T10:36:46-0400

Equation is separable:

"p^2-2px=2qy-q^2-1"

it follows that

"p^2-2px=a,\ 2qy-q^2-1=a", for constant a

Solving the first equation for p

"p^2-2px-a=0"

using quadratic formula

"p_{12}=x\pm\sqrt{x^2+a}"

and the second equation

"q^2-2qy+1+a=0"

"q_{12}=y\pm\sqrt{y^2-1-a}"

from relation

"dz=pdx+qdy," where z is solution of the equation

"dz=(x\pm\sqrt{x^2+a})dx+(y\pm\sqrt{y^2-1-a})dy" (four possible combinations of signs + and -)

"z=\int(x\pm\sqrt{x^2+a})dx+\int(y\pm\sqrt{y^2-1-a})dy"

"z=\frac{x^2+y^2}{2}\pm\int(\sqrt{x^2+a})dx\pm\int(\sqrt{y^2-1-a})dy"

"\int(\sqrt{x^2+a})dx=\\1/2 (x \sqrt{(a + x^2)} + a\cdot log(x + \sqrt{a + x^2}))+C_1"

"\int(\sqrt{y^2-1-a})dy=\\1/2(y\cdot\sqrt{-1 - a + y^2} - (1 + a) log(y + \sqrt{-1 - a + y^2}))+C_2"

"z=\frac{x^2+y^2}{2}\pm\\
1/2 (x \sqrt{(a + x^2)} + a\cdot log(x + \sqrt{a + x^2}))\pm\\
1/2(y\cdot\sqrt{-1 - a + y^2} - (1 + a) log(y + \sqrt{-1 - a + y^2}))+b,"

 a and b are real constants.

Answer:

"z=\frac{x^2+y^2}{2}\pm\\
1/2 (x \sqrt{(a + x^2)} + a\cdot log(x + \sqrt{a + x^2}))\pm\\
1/2(y\cdot\sqrt{-1 - a + y^2} - (1 + a) log(y + \sqrt{-1 - a + y^2}))+b,"

a and b are real constants.